OpenStudy (anonymous):

The sales at a particular bookstore grew from $2013 million in 2000 to $3864 million in 2005. Find an exponential function to model the sales as a function of years since 2000. Give your answer using the form B = Bo *a^t. a.) Bo= millions b.) a = (exact value) Please help!! :)

OpenStudy (anonymous):

will year 2000 be considered t =0? or t=2000?

OpenStudy (anonymous):

I'm not sure I think its 0

OpenStudy (anonymous):

ok then if at year 2000, t=0 i will be writing it in the form of a function B(t) = Bo *a^t -- where Bo and a are constants we know that B(0)=2013= Bo*a^0 since a^0 = 1, you can quickly solve for Bo now that you have Bo, solve for a we know that B(5)=3864=Bo*a^5 so just solve for a

OpenStudy (anonymous):

if you're interested in solving it when t=2000 then ill also explain that too you it is much harder than t=0

OpenStudy (anonymous):

you can try that too, just in case.

OpenStudy (anonymous):

alright we'll start with the the given function \[B=B_oa^t\] find the natural log of both sides \[ln B=ln ({B_oa^t})\] using the properties of logs \[ln B=ln {B_o}+ln ({a^t})\] still using the properties of logs \[ln B=ln {B_o}+t*ln ({a})\] then when you substitute in the actual values, you will get a system of equations 2 equations 2 unknowns. you just solve that in which ever method you know best B=2013 when t=2000 B=3864 when t=2005 \[ln (2013)=ln {B_o}+2000*ln (a)\] \[ln (2864)=ln {B_o}+2005*ln ({a})\] if it helps, you can substitute x=ln(B_o) y=ln(a) then once you find x and y B_o=e^x a=e^y

OpenStudy (anonymous):

ln(2013)=x+2000y ln(2864)=x+2005y

OpenStudy (anonymous):

I'm not getting what a is. I got Bo^t is which is 2013 millions.

OpenStudy (anonymous):

first method or 2nd method?

OpenStudy (anonymous):

first.

OpenStudy (anonymous):

Can you just tell me the answer so I can figure it out from there? Please.

OpenStudy (anonymous):

3864=2013*a^5 3864/2013=a^5 (3864/2013)^(1/5)=(a^5)^(1/5) x^(1/5) is the same as the 5th root a=(3864/2013)^(1/5)

OpenStudy (anonymous):

Okay I got it. Thank you!!

OpenStudy (anonymous):

you might wanna help out your friend